I need help solving the rest of this math question. May answer for the circumference is 62.83 ft, but I didn't know how to calculate the area.
What is the circumference and area in feet of circular pool that has a diameter of 20 yards?

Answer by ankor@dixie-net.com(22740)

What is the circumference and area in feet of circular pool that has a diameter of 20 yards?
:
You should be familiar with the "Area of a circle equation: A = pi*r^2
:
If the diameter is 20, the radius is 10,
:
A = pi*10^2
A = pi*100
A = 314.16 sq/ft